mayathegamergrl:

Does doubling the height of a cylinder have the same effect on the volume of a cylinder as doubling the radius? Please and thankss

kittybasil:

Infodump! Do you recall the volume of a cylinder? Here's the formula:\[V_{cylinder}=\pi r^2h\]So what's the difference between doubling the height:\[V_1=\pi r^2\color{red}{2h}\]and doubling the radius:\[V_2=\pi(\color{red}{2r})^2h\]


Leaving the rest up to Ex of course. He got here first

mayathegamergrl:

okay, thankss

kittybasil:

no, give him the medal, stopit

kittybasil:

Wait where'd he go

mayathegamergrl:

1 attachment
Extrinix:

(Make sure you read though it to understand it) Ok, Mexy, here is an example just to help make it clear: \(V_\text{cylinder}=\pi r^2 h\) now, take this and add in simple numbers to the variables \(V_\text{cylinder}=\pi (10)^2 (5)\) convert it into the doubling of the height: \(V_\text{x2height}=\pi (10)^2 (2(5))\) which gives you `3141.6` convert it into the doubling of the radius: \(V_\text{x2radius}=\pi (2(10))^2 (5)\) which gives you `6283.2` This means that, `no`, the doubling of the height is the same as the doubling of the radius.

mayathegamergrl:

alr tysm ext

kittybasil:

Medal for the man of the hour!! (🏅rain ensues)

Extrinix:

You're welcome, and don't forget to thank Kitt for that data dump, that rascal. .-.

kittybasil:

extrinix wrote:
You're welcome, and don't forget to thank Kitt for that data dump, that rascal. .-.
But of course...

1 attachment
12jacklopez:

ok